-1/6m - 5/8 + 3/4m = 1 3/4
M= ?

Answer by ewatrrr(24785)

-1/6m - 5/8 + 3/4m = 1 3/4
-1/6m - 5/8 + 3/4m = 7/4 LCD = 24
(-4/24)m - 15/24+ (18/24)m = 42/24
(14/24)m = 42/24 + 15/24
(14/24)m = 57/24
m = (57/24)(24/14)
m = 57/14 0r 4 1/14